Suppose Komla throws a die repeatedly until he gets a six. What is the probability
that he needs to throw more than 10 times to get a six, to 3 decimal places
1
Expert's answer
2020-06-01T19:30:28-0400

The event we are after is equivalent to "no six on the first 10 rolls". Since this requires any number other than 6 on the first roll (p=5/6), any number other than 6 on the second roll, etc the probability is

P=(5/6)10=0.162